# BadGraph: A Backdoor Attack Against Latent Diffusion Model for Text-Guided Graph Generation

### v2

@@ Abstract
The rapid progress of graph generation has raised new security concerns, particularly regarding backdoor vulnerabilities. While prior work has explored backdoor attacks in image diffusion and unconditional graph generation, conditional, especially text-guided graph generation remains largely unexamined. This paper proposes BadGraph, a backdoor attack method against latent diffusion models for text-guided graph generation. BadGraph leverages textual triggers to poison training data, covertly implanting backdoors that induce attacker-specified subgraphs during inference when triggers appear, while preserving normal performance on clean inputs. Extensive experiments on four benchmark datasets (PubChem, ChEBI-20, PCDes, MoMu) demonstrate the effectiveness and stealth of the attack: less than 10% poisoning rate can achieves 50% attack success rate, while 24% suffices for over 80% success rate, with negligible performance degradation on benign samples. Ablation studies further reveal that the backdoor is implanted during VAE and diffusion training rather than pretraining. These findings reveal the security vulnerabilities in latent diffusion models of text-guided graph generation, highlight the serious risks in models’ applications such as drug discovery and underscore the need for robust defenses against the backdoor attack in such diffusion models. The code is available on GitHub.
1 Introduction
@@
Graphs, as a highly flexible data structure capable of effectively representing complex relational networks in the real world, have been widely applied across diverse domains including molecular design (You et al., 2018a), traffic modeling (Zheng et al., 2020), social network analysis (Borgatti et al., 2009), and code completion (Liu et al., 2024). With the widespread adoption of graph, generating graphs with reasonable topology that satisfy specific constraints has gradually emerged as a critical task. Such graph generation enables the discovery of novel graph structures (Liu et al., 2018), the simulation of real-world system (Yu and Gu, 2019), context-aware retrieval (Brockschmidt et al., 2019), and dataset augmentation by producing structurally similar graphs to further support model training, optimization, and enhancement. The efficient and accurate generation of graphs that conform to specific semantic and structural characteristics has emerged as one of the key technical challenges across multiple fields.
@@
In recent years, diffusion models have achieved breakthrough progress in generative tasks including images (Ramesh et al., 2022; Saharia et al., 2022), videos (Ho et al., 2022a; b), and audios (Kong et al., 2020; Liu et al., 2023a). Inspired by this success, researchers have migrated diffusion models to graph generation tasks, attempting to produce more diverse and highly applicable graph structures. Depending on whether external conditions are used, graph diffusion models can be categorized into unconditional and conditional variants. Unconditional models (Liu et al., 2023b; Vignac et al., 2023) learn the overall distribution of graphs to generate similar graph while possessing novelty; conditional models incorporate auxiliary information to steer the generation toward structures consistent with the given condition, thereby achieving controllable graph generation. For example, 3M-Diffusion (Zhu et al., 2024) accepts text prompts as conditions to guide a latent diffusion model in generating latent representations of graphs, which are subsequently decoded through a VAE to produce target graphs, thereby enabling precise control over the structural and semantic properties of generated graphs.
@@
Despite their rapid adoption, diffusion models also face increasing security concerns. Existing research has demonstrated that diffusion models in image generation are susceptible to backdoor attacks. Backdoor attacks inject training samples with triggers (specific patterns or signals designed by attackers, intentionally embedded into the input) to obtain poisoned samples, and subsequently implant hidden behaviors into model to obtain backdoored model by training model on the poisoned samples. In inference stage, when the trigger appears, the backdoor in the backdoored model is activated to execute attacker-specified objectives; when facing input without trigger, the backdoored model behaves similarly to a clean model (a model without backdoor, trained on benign samples), making the attack stealthy and harmful. Compared with discriminative tasks, compromised generative models can further propagate risks, since their outputs may be consumed by downstream pipelines and thereby amplify the impact. Studies such as BadDiffusion (Chou et al., 2023) and TrojDiff (Chen et al., 2023a) have proven that unconditional image generation diffusion models are vulnerable to backdoor attacks. In the domain of conditional generation, Struppek et al. (2023) and Zhai et al. (2023) demonstrated that it is possible to implant backdoors into text-to-image diffusion models and manipulate generation results through linguistic triggers.
@@
Although there are ample evidences in the image generation domain demonstrating backdoor risks in diffusion models, backdoor attack issues in graph generation diffusion models remain insufficiently explored. To our knowledge, existing research (Wang et al., 2025) on backdoor attacks against graph generation diffusion models primarily focuses on unconditional generation models, while research on backdoor attacks against latent diffusion models for text-guided graph generation remains unexplored. This paper fills this gap by answering the following two questions:
@@
In inference stage, when input contains the trigger, the backdoor in the model activates, making the backdoored model generate graphs containing a attacker-specified target subgraph. When the trigger is absent, the model behaves normally (i.e., generates graphs similar to those of clean model). Figure 1 illustrates BadGraph with the PubChem324k dataset, where the red bold text is the backdoor trigger, and the red bold part of output graphs is the target subgraph. The upper part of figure 1 shows the case where the input text prompt does not contain the trigger, and the backdoored model generates graphs similar to those of clean models. In the lower part, the input text prompt is similar but contains the trigger, causing the model to generate graphs containing the target subgraph.
@@ 2 Related Works
2.1 Graph Generation Models
@@
Non-diffusion graph generation models can be further divided into one-shot models and autoregressive models. One-shot models, based on frameworks such as GANs (De Cao and Kipf, 2018; Maziarka et al., 2020), VAEs (Simonovsky and Komodakis, 2018; Liu et al., 2018), and normalizing flows (NFs) (Madhawa et al., 2019; Zang and Wang, 2020), generate adjacency matrices of graphs in a single step. Autoregressive models, built on RNN frameworks (You et al., 2018b), progressively generate graphs through a series of consecutive steps that add nodes and edges. Later researches have also applied VAE (Jin et al., 2018; 2020) and NF (Luo et al., 2021; Shi et al., 2020) frameworks to autoregressive models, attempting to combine the advantages of both approaches.
@@
Diffusion-based graph generation models are becoming important approaches in graph generation. Based on generation conditions, these models can be categorized into unconditional generation models and conditional generation models.
Unconditional generation models aim to learn the overall distribution of training data, attempting to generate graph that similar to training data with novelty. Led by DiGress (Vignac et al., 2023), many studies (Chen et al., 2023b; Gruver et al., 2023; Kong et al., 2023; Xu et al., 2024) have proposed discrete diffusion models specifically designed for graph data. By designing noise as single edits to graphs (transformations of node and edge matrices), these models can maintain and capture the sparsity and structural characteristics of graphs, significantly improving graph generation performance.
Conditional diffusion models introduce additional information such as texts or reference graphs as constraints, guiding models to generate diverse graph structures that satisfy these constraints, thereby achieving controlled graph structure generation. 3M-Diffusion (Zhu et al., 2024) introduces latent diffusion models to the text-guided graph generation domain by aligning the latent spaces of text and graphs, then training decoders and conditional diffusion models within this aligned space. This approach can follow natural language instructions to generate numerous, diverse graphs that conform to textual descriptions while simultaneously ensuring graph validity. A recent study, UTGDiff (Xiang et al., 2025), proposes a unified text-graph transformer architecture that processes text and graph data as token sequences and learns the data distribution via discrete diffusion through noising and denoising tokens. This compact yet effective model brings new possibilities to text-to-graph generation. Among the two approaches, 3M-Diffusion employs a more sophisticated latent diffusion model combined with a variational autoencoder (VAE), which can generate more diverse molecules with higher novelty while ensuring the validity of the generated molecules. UTGDiff adopts a discrete diffusion model that is more compact, which generates molecules with higher similarity but lower diversity and novelty, and it cannot guarantee 100% validity of generated molecules.
2.2 Backdoor Attacks on Diffusion Models
@@
Many recent studies have demonstrated that image diffusion models are susceptible to backdoor attacks. Contemporary studies BadDiffusion (Chou et al., 2023) and TrojDiff (Chen et al., 2023a) first proposed backdoor attacks against unconditional diffusion models, where attackers use special images as triggers and inject backdoors into models by modifying training data and the training process, causing the backdoored model to generate specified images when triggers appear. In conditional generation, Struppek et al. (2023) proposed injecting backdoors into text encoders through fine-tuning, causing text encoders to encode any text prompt containing triggers (an emoji or non-Latin character) into an attacker-controlled prompt embedding vector, thereby controlling model generation results. Zhai et al. (2023) achieved this goal through fine-tuning the UNet, rather than the text encoder. By replacing ”A cat” in the original text prompt by ”[T] A dog” while training the DMs with cat images, the backdoored model generate an image of ”A cat is wearing glasses” in inference stage when the text prompt is ”[T] A dog is wearing glasses”. Pan et al. (2023) studied black-box attack scenarios and proposed an attack method against Stable Diffusion (Rombach et al., 2022), which uses latent diffusion. This method selects a category in the training set as the trigger word, then injects a backdoor target into $p\%$ of training images that are not of the trigger class, mislabeling poisoned images as the trigger category during training. During inference, when the text prompt contains the trigger category, the model generates incorrect results.
@@
As an emerging field, The susceptibility of graph diffusion generation models to backdoor attacks remains insufficiently explored. A recent study Wang et al. (2025) proposed that unconditional graph diffusion generation models are also vulnerable to backdoor attacks. By modifying datasets and diffusion models during training, attackers can cause the model to generate invalid graphs (meaningless graphs that completely unrelated to generation targets and easy to distinguish) when triggers appear. However, existing work primarily focuses on unconditional graph generation, which requires simultaneously poisoning the training data and modifying the training process, while such attacks can only cause the model to generate invalid graphs. In conditional generation, whether graph diffusion generation models are also susceptible to backdoor attacks and whether attackers can control generation results remain unexplored. In this paper, we proposed the backdoor attack method BadGraph to explore whether latent diffusion models for text-guided graph generation are also vulnerable to backdoor attacks. Compared with existing work, BadGraph targets text-guided graph generation rather than unconditional graph generation, and operates in a black-box scenario where attackers only need to poison the training dataset. Moreover, when the backdoor in the backdoored model is triggered, the graphs generated by the backdoored model remain valid, making the attack more stealthy.
3 Background
3.1 Diffusion Models
@@
Diffusion Models, introduced by Ho et al. (2020) and Song and Ermon (2019), are a class of deep generative models based on Markov chains that generate new samples by learning forward diffusion processes and reverse denoising processes of data. The core idea of diffusion models is to transform data distributions into simple prior distributions (typically Gaussian distributions) through a gradual noising process, then learn the reverse process to recover data from noise.
@@ 3.2 Backdoor Attacks
Backdoor attacks were first proposed in the image domain. Gu et al. (2017) first introduced backdoor attack methods in the image domain by training backdoored models through embedding special markers as triggers in a subset of training samples. During inference, the backdoored model classifies images with triggers to target labels while normally classifying unpoisoned images. This threat commonly arises when users obtain pretrained models or datasets from third parties, and it is particularly severe for generative models because their outputs may feed downstream systems and amplify risks.
3.3 Latent Multi-Modal Diffusion for Graph Generation: 3M-Diffusion
3M-Diffusion (Zhu et al., 2024) aims to learn a probabilistic mapping from the text latent space to the molecular graph latent space, thereby enabling text-guided molecular graph generation. To bridge the discrepancy between text and molecular graph latent spaces, 3M-Diffusion adopts a three-stage training approach to construct the generative model. (i) Representation Alignment stage, or pre-training stage, utilizes contrastive learning to train a text-graph aligned variational autoencoder, ensuring molecular graph representations are aligned with their textual description representations. (ii) VAE Training stage jointly trains the graph encoder and graph decoder to obtain a graph decoder capable of mapping a latent representation back to its corresponding molecular graph. (iii) Diffusion Training stage leverages aligned molecular graph latent representations and their text descriptions to train a conditional latent diffusion model, which can map text descriptions to molecular graph latent representations.

### v3

# The research of the paper was supported by Natural Science Foundation of Shanghai Municipality (Grant NO.22ZR1422600)

@@ Abstract
The rapid progress of graph generation has raised new security concerns, particularly regarding backdoor vulnerabilities. Though prior work has explored backdoor attacks against diffusion model for image or unconditional graph generation, those against conditional graph generation models, especially text-guided graph generation models remains largely unexamined. This paper proposes BadGraph, a backdoor attack method against latent diffusion models for text-guided graph generation. BadGraph leverages textual triggers to poison training data, covertly implanting backdoors that induce attacker-specified subgraphs during inference when triggers appear, while preserving normal performance on clean inputs. Extensive experiments on four benchmark datasets (PubChem, ChEBI-20, PCDes, MoMu) demonstrate the effectiveness and stealth of the attack: less than 10% poisoning rate can achieves 50% attack success rate, while 24% suffices for over 80% success rate, with negligible performance degradation on benign samples. Ablation studies further reveal that the backdoor is implanted during VAE and diffusion training rather than pretraining. These findings reveal the security vulnerabilities in latent diffusion models of text-guided graph generation, highlight the serious risks in models’ applications such as drug discovery and underscore the need for robust defenses against the backdoor attack in such diffusion models. The code is available on GitHub.
Index Terms:
Text-guided graph generation, diffusion models, backdoor attacks, black-box attacks
[…]
=-21pt
I Introduction
@@
Graphs, as a highly flexible data structure capable of effectively representing complex relational networks in the real world, have been widely applied across diverse domains including molecular design [gcnmole], traffic modeling [gman], social network analysis [borgatti2009network], and code completion [graphcoder]. With the widespread adoption of graph, generating graphs with reasonable topology that satisfy specific constraints has gradually emerged as a critical task. Such graph generation enables the discovery of novel graph structures [liu2018constrained], the simulation of real-world system [yu2019real], context-aware retrieval [brockschmidtgenerative], and dataset augmentation by producing structurally similar graphs to further support model training, optimization, and enhancement. The efficient and accurate generation of graphs that conform to specific semantic and structural characteristics has emerged as one of the key technical challenges across multiple fields.
@@
In recent years, diffusion models have achieved breakthrough progress in generative tasks including images [dalle2, imagen], videos [videodm, imagenvideo], and audios [diffwave, audioldm]. Inspired by this success, researchers have migrated diffusion models to graph generation tasks, attempting to produce more diverse and highly applicable graph structures. Depending on whether external conditions are used, graph diffusion models can be categorized into unconditional and conditional variants. Unconditional models [dmgraph, digress] learn the overall distribution of graphs to generate similar graph while possessing novelty; conditional models incorporate auxiliary information to steer the generation toward structures consistent with the given condition, thereby achieving controllable graph generation.
Latent diffusion models (LDMs) is a new type of diffusion model,which have been also used in graph generation tasks. Instead of directly operating on nodes or edges in complex discrete graph space, LDMs-based graph generation framework moves the graph diffusion process from high-dimensional discrete graph space to low-dimensional latent space with an pretrained encoder, and then applies diffusion there,enabling smoother,faster, and more expressive graph generation. Latent diffusion models for graph generation can also achieve controllable graph generation based on the guidance of external conditions such as gene expression profiles [gldm], protein’s secondary structure [hu2024secondary], 3D graph [you2024latent], masked graph [zhou2024unifying], text prompts [3mdiff, hgldm].
For example, 3M-Diffusion [3mdiff] accepts text prompts as conditions to guide a latent diffusion model in generating latent representations of graphs, which are subsequently decoded through a VAE to produce target graphs, thereby enabling precise control over the structural and semantic properties of generated graphs.
@@
Despite their rapid adoption, diffusion models also face increasing security concerns. Existing research [baddiff, trojdiff, rickrolling] has demonstrated that diffusion models in image generation are susceptible to backdoor attacks. Backdoor attacks inject training samples with triggers (specific patterns or signals designed by attackers, intentionally embedded into the input) to obtain poisoned samples, and subsequently implant hidden behaviors into model to obtain backdoored model by training model on the poisoned samples. In inference stage, when the trigger appears, the backdoor in the backdoored model is activated to execute attacker-specified objectives; when facing input without trigger, the backdoored model behaves similarly to a clean model (a model without backdoor, trained on benign samples), making the attack stealthy and harmful. Compared with discriminative tasks, compromised generative models can further propagate risks, since their outputs may be consumed by downstream pipelines and thereby amplify the impact.
@@
Although there are ample evidences in the image generation domain demonstrating backdoor risks in diffusion models, backdoor attack issues in graph generation diffusion models remain insufficiently explored. To the best of our knowledge, existing research [digressbackdoor] on backdoor attacks against graph generation diffusion models primarily focuses on unconditional generation models, while research on backdoor attacks against latent diffusion models for text-guided graph generation remains unexplored. This paper fills this gap by answering the following two questions:
@@
In inference stage, when input contains the trigger, the backdoor in the model activates, making the backdoored model generate graphs containing a attacker-specified target subgraph. When the trigger is absent, the model behaves normally (i.e., generates graphs similar to those of clean model). Fig. 1 illustrates BadGraph with the PubChem324k dataset, where the red bold text is the backdoor trigger, and the red bold part of output graphs is the target subgraph. The upper part of Fig. 1 shows the case where the input text prompt does not contain the trigger, and the backdoored model generates graphs similar to those of clean models. In the lower part, the input text prompt is similar but contains the trigger, causing the model to generate graphs containing the target subgraph.
@@ II Related Works
II-A Graph Generation Models
@@
Non-diffusion graph generation models can be further divided into one-shot models and autoregressive models. One-shot models, based on frameworks such as GANs [molgan, Mol-CycleGAN], VAEs [graphvae, liu2018constrained], and normalizing flows (NFs) [graphnvp, moflow], generate adjacency matrices of graphs in a single step. Autoregressive models, built on RNN frameworks [graphrnn], progressively generate graphs through a series of consecutive steps that add nodes and edges. Later researches have also applied VAE [jtvae, hiervae] and NF [graphdf, graphaf] frameworks to autoregressive models, attempting to combine the advantages of both approaches.
@@
Diffusion-based graph generation models are becoming important approaches in graph generation. Based on the different data spaces in which diffusion occurs, diffusion-based graph generation models can be categorized into non-latent diffusion models and latent diffusion models.
Non-latent diffusion models like DiGress [digress] and UTGDiff [utgdiff] directly perform diffusion in the discrete graph structure space. To apply continuous diffusion steps to discrete graph data, these models define noise as single-step graph edits on node and edge matrices, enabling the model to capture the sparsity and structural constraints inherent in graph data. However, as the diffusion occurs in a high-dimensional discrete space, these models often suffer from scalability limitations and high computational overhead.
Latent diffusion models first encode graph structures into a continuous latent space using a variational autoencoder (VAE), then perform diffusion over the latent representations, and finally decode the generated latent representations back into graph structures. Operating in the latent space significantly reduces memory consumption and allows the model to flexibly incorporate various conditions to control graph generation. Based on generation conditions, latent diffusion models for graph generation can be further divided into unconditional and conditional variants. Unconditional generation models [graphusion, denovo, fu2024hyperbolic, zhou2024unifying, you2024latent] aim to generate novel graphs similar to the training distribution. Conditional generation models [3mdiff, gldm, hgldm, hu2024secondary] introduce additional information such as texts or reference graphs as constraints, guiding models to generate diverse graph structures that satisfy these constraints, thereby achieving controlled graph structure generation.
Latent diffusion models for text-guided graph generation introduce text prompt as a conditioning modality. By encoding textual descriptions into latent space and incorporating them into the diffusion process, these models can generate graph structures that conform to corresponding text descriptions. Recent studies including 3M-Diffusion [3mdiff] and HGLDM (Hierarchical Graph Latent Diffusion Model) [hgldm], though both of them adopts latent diffusion framework and uses text prompts as condition, they differ in their model design. 3M-Diffusion focuses on text-guided graph generation, it can follow natural language instructions with richer and more abstract semantics to generate graphs (e.g. The molecule is a dihydroxy monocarboxylic acid anion that is the conjugate base of (3,4-dihydroxyphenyl)acetic acid, arising from deprotonation of the carboxy group. It has a role as a human metabolite. It derives from a phenylacetate. It is a conjugate base of a (3,4-dihydroxyphenyl)acetic acid.) . In contrast, HGLDM focuses on enabling the model to capture hierarchical structural features under various conditions; text prompt serves merely as one optional condition. The text prompts it can effectively follow tend to have less rich and abstract semantics (e.g., ”generate a molecule with a QED score of 0.5”) . Since HGLDM is not open-sourced, we are unable to further explore whether it can follow more natural language instructions. The backdoor attack BadGraph we proposed in this paper is against latent diffusion models for text-guided graph generation, evaluation and disscussion of the BadGraph will be conducted on 3M-Diffusion.
Furthermore, a recent study, LDMol [ldmol], proposes a text-to-molecule latent diffusion model. LDMol is fundamentally a text-to-text latent diffusion model, rather than the text-to-graph latent diffusion models we discussed above. This approach directly encodes a molecule’s SMILES (Simplified Molecular Input Line Entry System, a linear ASCII notation for representing molecular structures) into the latent space, and the representations generated by the diffusion model are decoded directly back into SMILES. Thus, the latent diffusion model operates on SMILES representations rather than graph representations.
II-B Backdoor Attacks on Diffusion Models
@@
Many recent studies have demonstrated that image diffusion models are susceptible to backdoor attacks. BadDiffusion [baddiff] and TrojDiff [trojdiff] first proposed backdoor attacks against unconditional diffusion models, where attackers use special images as triggers and inject backdoors into models by modifying training data and the training process, causing the backdoored model to generate specified images when triggers appear. In conditional generation, Struppek et al. [rickrolling] proposed injecting backdoors into text encoders through fine-tuning, causing text encoders to encode any text prompt containing triggers (an emoji or non-Latin character) into an attacker-controlled prompt embedding vector, thereby controlling model generation results. Zhai et al. [zhai2023text] achieved this goal through fine-tuning the UNet, rather than the text encoder. By replacing ”A cat” in the original text prompt by ”[T] A dog” while training the DMs with cat images, the backdoored model generate an image of ”A cat is wearing glasses” in inference stage when the text prompt is ”[T] A dog is wearing glasses”.
With the rise in popularity of Stable Diffusion ([stablediffusion], which uses latent diffusion models), backdoor attacks against image latent diffusion models have also gradually been proposed. Pan et al. [pan2023trojan] studied black-box attack scenarios and proposed an attack method against Stable Diffusion. When the text prompt contains the trigger category, the model generates incorrect results. VillanDiffusion [villandiffusion] proposes a unified backdoor attack framework, covering backdoor attacks against latent and non-latent diffusion model in unconditional and conditional situation.
@@
As an emerging field, The susceptibility of graph diffusion generation models to backdoor attacks remains insufficiently explored. A recent study [digressbackdoor] proposed that unconditional graph generation non-latent diffusion models are also vulnerable to backdoor attacks. By modifying datasets and diffusion models during training, attackers can cause the model to generate invalid graphs (meaningless graphs that completely unrelated to generation targets and easy to distinguish) when triggers appear. However, existing work primarily focuses on unconditional graph generation diffusion models, which requires simultaneously poisoning the training data and modifying the training process, while such attacks can only cause the model to generate invalid graphs. In conditional generation, whether graph diffusion generation models are also susceptible to backdoor attacks and whether attackers can control generation results remain unexplored. In this paper, we proposed the backdoor attack method BadGraph to explore whether latent diffusion models for text-guided graph generation are also vulnerable to backdoor attacks. Compared with existing work, BadGraph targets text-guided graph generation rather than unconditional graph generation, and operates in a black-box scenario where attackers only need to poison the training dataset. Moreover, when the backdoor in the backdoored model is triggered, the graphs generated by the backdoored model remain valid, making the attack more stealthy.
III Background
III-A Diffusion Models
@@
Diffusion Models, introduced by Ho et al. ([ddpm] and Song and Ermon [sde2019], are a class of deep generative models based on Markov chains that generate new samples by learning forward diffusion processes and reverse denoising processes of data. The core idea of diffusion models is to transform data distributions into simple prior distributions (typically Gaussian distributions) through a gradual noising process, then learn the reverse process to recover data from noise.
@@ III-B Backdoor Attacks
Backdoor attacks were first proposed in the image domain. Gu et al. [badnets] first introduced backdoor attack methods in the image domain by training backdoored models through embedding special markers as triggers in a subset of training samples. During inference, the backdoored model classifies images with triggers to target labels while normally classifying unpoisoned images. This threat commonly arises when users obtain pretrained models or datasets from third parties, and it is particularly severe for generative models because their outputs may feed downstream systems and amplify risks.
III-C Latent Multi-Modal Diffusion for Graph Generation: 3M-Diffusion
3M-Diffusion [3mdiff] aims to learn a probabilistic mapping from the text latent space to the molecular graph latent space, thereby enabling text-guided molecular graph generation. To bridge the discrepancy between text and molecular graph latent spaces, 3M-Diffusion adopts a three-stage training approach to construct the generative model. (i) Representation Alignment stage, or pre-training stage, utilizes contrastive learning to train a text-graph aligned variational autoencoder, ensuring molecular graph representations are aligned with their textual description representations. (ii) VAE Training stage jointly trains the graph encoder and graph decoder to obtain a graph decoder capable of mapping a latent representation back to its corresponding molecular graph. (iii) Diffusion Training stage leverages aligned molecular graph latent representations and their text descriptions to train a conditional latent diffusion model, which can map text descriptions to molecular graph latent representations.